OpenStudy (anonymous):

at decreasing and increasing function what is meaning for undefined and when the function has open interval and close interval ?

OpenStudy (mathmagician):

lets take \[f(x)=\left(\begin{matrix}x, x \in (- \infty,1)\\ -x, x \in (1,+\infty)\end{matrix}\right)\]. This function is defined in open intervals \[(- \infty, 1)\] and \[(1,+ \infty)\], but undefines at point x=1